What isthe horse supplied by the driving mootor to pum tack 1L ~1kg if height is 50meter calculate Hp?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
 
Work done by motor per sec W = mgh = 1*10*50 = 500J
1HP = 746J/s
then HP supplied to driving motor = 500/746 = 0.67HP
 
 
 
Regards
Arun (askIITians forum expert)